The resistance (in ohms) of a circular conductor varies directly with the length of the conductor and inversely with the square of the radius of the conduction. If 50ft of wire with a radius of 6 x 10^-3 inch has a resistance of 10ohms, what would be the resistance of 100ft of the same wire if the radius is increased to 7 x 10^13 inch?
Answer by lwsshak3(11628)

he resistance (in ohms) of a circular conductor varies directly with the length of the conductor and inversely with the square of the radius of the conduction. If 50ft of wire with a radius of 6 x 10^-3 inch has a resistance of 10ohms, what would be the resistance of 100ft of the same wire if the radius is increased to 7 x 10^13 inch?
**
Resistance=10(100/50)[(6*10^-3)/(7*10^13)]^2
=20(6*10^-16)/7)=(120/7)*10^-16≈17.14*10^-16 ohms